In a certain electrolysis experiment involving Al3+ ions, 60.2 kg of Al is


recovered when a current of 0.325 A is used. How many minutes did the


electrolysis last?

1
Expert's answer
2021-12-27T12:13:02-0500

Atomic weight of Al = 27

Reaction: Al⁺³ + 3e = Al

3 moles of electron is required by 1 mole of Al

Proportion:

(3 moles×96500 C/mol) of charge – for 27 g of Al

x of charge – for 60.2 g of Al

x = 645.47×10³ C – charge required

Charge = Current×time

Time = charge/current = (645.47×10³ C) / (0.352 A) =1833×10³ s = 30.55×10³ min

Answer: 30.55×10³ min